INVESTIGATION OF O_{2}(C^{3}Π, v=2) BY NOVEL LASER PHOTOIONIZATION TECHNIQUE IN AIR AT ATMOSPHERIC PRESSURE

\underline{JONATHAN D. UMBEL}, STEVEN F. ADAMS, CHARLES A. DeJOSEPH, JR., Propulsion Directorate, Air Force Research Laboratory, WPAFB, OH 45433.

We have used pulsed laser photoionization of dry atmospheric air for the analysis of the O_{2}(X^{3}Σ→C^{3}Π) excitation spectrum and improved characterization of the O_{2}(C^{3}Π) Ryberg state. This ^{3}Π Ryberg state is generally very diffuse due to predissociation, but using this method, reasonable resolution in the rotational structure of the v=2 level was achieved. A simulated O_{2}(X^{3}Σ→C^{3}Π) excitation spectrum was calculated for line position, intensity, and linewidth that agreed very well with the experiment, providing values for ν_{0}, B_{eff}, and D_{eff} for the F1, F2, and F3 subbands of O_{2}(C^{3}Π, v=2). The photoionization technique involved a combination of resonant-enhanced multi-photon ionization (REMPI), collisional excitation, and laser-induced fluorescence. A focused, ultraviolet laser pulse was used to excite the O_{2}(X^{3}Σ→C^{3}Π) resonant transition. The near coincident energy between the O_{2}(C^{3}Π, v=2) state and the N_{2}(a^{′1}Σ_{u}^{−}, v=1) state resulted in rapid collisional transfer of energy to the N_{2}(a′) state. The N_{2}(a′) molecules were then photoionized by the same laser pulse to the N_{2}^{+}(B) state, producing an easily observable fluorescence at 391 nm. By scanning the laser wavelength while observing the 391 nm fluoresence, details of the O_{2}(C^{3}Π) Ryberg state could be studied.